OpenStudy (anonymous):

solve for x: 2/5 + 3/5x = x+5/10 the answers are: x=-3 x=2 x=-3 and x=2 no solution please i need help :/

OpenStudy (anonymous):

Is that \[3/(5x)\] or \[(3/5) \times x\] also: \[(x + 5) / 10\] or \[x + (5/10)\] ??

OpenStudy (jlastino):

X=-3 X=2

OpenStudy (anonymous):

3/(5x)

OpenStudy (anonymous):

-0.25

OpenStudy (anonymous):

Well assuming you meant x + (5/10) on the right side: You're trying to isolate for x: (basically BEDMAS in reverse) 2/5 + 3/(5x) = x + 5/10 subtract 2/5 from each side: 3/(5x) = x + 5/10 - 2/5 subtract x from each side: 3/(5x) - x = 5/10 - 2/5 make a common denominator: 3/(5x) - x = 5/10 - 4/10 3/(5x) - x = 9/10 make another common denominator: (3 - 5x)/(5x) = 9/10 cross-multiply: (3 - 5x)(10) = (5x)(9) expand: 30 - 50x = 45x add 50x to both sides: 30 = 95x divide both sides by 95: x = 30/95 lol I got kinda carried away but I wanted to be verbose because I don't know your ability.